OpenStudy (anonymous):

What are the possible numbers of positive real, negative real, and complex zeros of f(x) = -7x^4 - 12x^3 + 9x^2 - 17x + 3? Positive Real: 3 or 1 Negative Real: 1 Complex: 2 or 0 Positive Real: 3 or 1 Negative Real: 2 or 0 Complex: 1 Positive Real: 1 Negative Real: 3 or 1 Complex: 2 or 0 Positive Real: 4, 2 or 0 Negative Real: 1 Complex: 0 or 1 or 3

OpenStudy (amorfide):

you want to substitute values into f(x) to start for example f(1) f(2) f(-1) etc until you find a value or two values that make f(x)=0 can you do this?

OpenStudy (anonymous):

okay let me try and sorry for replying late my internet keeps going off

OpenStudy (anonymous):

but what do i plug in for f(x)?

OpenStudy (amorfide):

f(x)= -7x^4 - 12x^3 + 9x^2 - 17x + 3 so f(1) means replace all x terms with 1

OpenStudy (anonymous):

okay so just to get this clear you want me to solve doing f(1) and f(2) and f(-1)

OpenStudy (amorfide):

i want you to use the value the question has given you Positive Real: 3 or 1 Positive Real: 4, 2 or 0

OpenStudy (anonymous):

okay so f(3) = -7x^4 - 12x^3 + 9x^2 - 17x + 3 f(3) = -7(3)^4 - 12(3)^3 + 9(3)^2 - 17(3) +3 f(3) = -7(81) - 12(27) + 9(9) - 17(3) +3 f(3) = - 567 - 324 + 81 - 51 + 3 f(3) = -858

OpenStudy (anonymous):

like that?

OpenStudy (amorfide):

yeah

OpenStudy (anonymous):

okay and if f(x) doesn't equal 0 its not the correct answer?

OpenStudy (amorfide):

oh ignore me

OpenStudy (anonymous):

what?

OpenStudy (amorfide):

i understood the question wrong

OpenStudy (anonymous):

oh

OpenStudy (amorfide):

okay so you know that the function is to the power of 4

OpenStudy (amorfide):

therefore it has 4 possible roots

OpenStudy (anonymous):

ok

OpenStudy (amorfide):

you want to figure out, how many real it can have (pos or neg) and how many complex it has

OpenStudy (anonymous):

idk how to do that? i missed a couple of days of school now she wants me to make it up over this spring break

OpenStudy (amorfide):

@ganeshie8 can you help him ? I can't do it

OpenStudy (anonymous):

thanks anyways